# Generate, but Verify: Reducing Hallucination in Vision-Language Models with Retrospective Resampling

### v2

@@ 1 Introduction
In this paper, we introduce REVERSE (REtrospective VERification and SElf-correction), the first framework that unifies generation adjustment with online post-hoc verification in a single VLM architecture (Figure 1). REVERSE consists of two key components. First, our method fine-tunes a VLM on a specially constructed training dataset consisting of synthetic hallucination phrases tagged by a special token. Unlike prior VLMs instructed with only well-grounded data, our resulting hallucination-aware model is now able to tag likely phrase-level hallucinations during the generation process.
Second, we introduce retrospective resampling, a technique that allows the hallucination-aware VLM to serve as its own verifier. During the generation process, when the hallucination-aware VLM places sufficient probability on the special hallucination token we trigger a backtracking self-correction process. Specifically, we backtrack to a previous confident section and then apply rejection sampling and query-rewriting to correct the hallucination. As illustrated in Figure 1, both components in REVERSE combine to effectively mitigate visual hallucinations.
@@ 4 Experiments
Implementation Details We applied our method, REVERSE, on three VLM backbones: LLaVA-v1.5 (7B) [34], LLaVA-More (LLaVA with Llama-v3.1 8B) [16, 1], and Qwen2.5-VL (3B) [6]. Since LLaVA provides both its pre-trained model and instruction tuning data, we performed LoRA fine-tuning on the pre-trained model directly with our modified cross-entropy loss (see subsection 3.2) and the 1.3M-sample dataset for one epoch. In contrast, Qwen2.5-VL does not release its instruction tuning data. To enable a fair comparison, we perform full fine-tuning on the publicly available Qwen2.5-VL model using two alternatives: a 100k subset of LLaVA’s instruction data and a matched subset from our dataset. Although a more direct evaluation using Qwen2.5-VL’s original instruction tuning data and the augmentation method described in subsection 3.1 would be ideal, it is not feasible given the current release conditions. More details on training recipes are provided in Appendix F.
@@
Evaluation Protocol To evaluate our approach, we compare REVERSE with various hallucination mitigation methods, including training-free or training-based generative adjustment techniques [27, 22, 15, 3, 53, 39, 54, 56], and post-hoc verification with refinement [51]. All methods are evaluated on both VLM backbones under consistent settings, where we fix the decoding temperature at 0 and use only the base prompts provided by each dataset to ensure fair comparisons. Since REVERSE does stochastic sampling at inference time, we report the mean performance over 100 bootstrapped runs for robustness. The exact numbers with 95% confidence intervals are provided in Appendix E.
Our evaluation dataset covers several standard VQA tasks aimed at assessing visual hallucination, with a primary focus on image captioning and open-ended question answering. While discriminative tasks, binary (Yes/No) questions targeting object, attributes, and spatial understanding, are also common, backtracking provides limited benefit in such settings, which have been noted to offer less diagnostic insight into VLM hallucinations [7, 39]. We include results on these tasks in Appendix E for completeness.
For image captioning, we use CHAIR-MSCOCO [38, 53] and the generative subset of AMBER [45] (denoted as AMBER-G). CHAIR-MSCOCO evaluates object hallucination using the CHAIR score, which measures the degree of misalignment between objects mentioned in a model-generated caption and objects actually present in the image. It is defined as 1 minus the intersection over union (IoU) between the sets of mentioned and ground-truth objects. We report both CHAIRi, which aggregates the CHAIR score across all object instances, and CHAIRs, which quantifies the proportion of images where at least one hallucination occurs. For AMBER-G, we report four key metrics: CHAIR, Coverage (Cover), Hallucination (Hall), and Cognition (Cog). CHAIR is the same as CHAIRi above, and Coverage measures how well the caption mentions all objects in the image, similar to recall. The definitions of the remaining two metrics and further details on these datasets are provided in Appendix D.
@@ 4.1 Experimental Results
Open-ended Question Answering We evaluate REVERSE on MMHal-Bench and HaloQuest, two open-ended VQA benchmarks containing lots of questions with false premises or insufficient context. These examples require the model to either refuse or correct the query. For these questions, we observed that REVERSE often produces empty responses and we interpret this behavior as the model identifying the query as unanswerable. In all such cases, we apply query rewriting with the prompt: “For this question, please point out the false premises or note what information is missing, rather than answering it directly.” The complete mechanism for handling unanswerable questions is given in Appendix F.
@@ 4.2 Discussions
We conduct additional experiments on two LLaVA models to examine key aspects of our method, including ablation studies, trade-offs between expressiveness and performance, efficiency versus accuracy, and the effect of temperature. Potential limitations and broader social impacts are discussed in Appendix C.
Ablation Studies We conduct ablation studies to evaluate the contributions of different components of our method, as shown in Table 4. Comparing the first and second rows, we observe that hallucination-aware training alone already improves performance across all metrics, outperforming existing VLMs. We hypothesize that this improvement arises from the model’s ability to contrast positive and negative phrases, effectively learning to distinguish between </CN> and </UN> during training—a mechanism that may be similar to DPO [37]. This finding suggests a potential research direction for future work. Interestingly, even a naive rejection sampling strategy reduces CHAIR hallucination scores by 1.2. When combined with query rewriting, the coverage improves by a further 1.2 points, indicating that rewriting helps the model explore alternative phrasing and correct itself more effectively.
Trade-offs Between Expressiveness and De-Hallucination As discussed in subsection 3.3, a key component of our retrospective resampling method is the predefined threshold . When the predicted probability of </UN> exceeds , the model triggers backtracking and self-correction. Figure 5 presents an analysis of the effect of on two VLMs. The 2D plot illustrates the performance trade-off between CHAIR (hallucination metric) and coverage across different threshold values. Based on our experiments, we selected as a global threshold for image captioning and other generative tasks, as it represents the peak of the performance frontier. Compared with prior hallucination reduction methods, REVERSE enables us to dynamically adjust the balance between expressiveness and trustworthiness. Notably, with a relatively high threshold of , our method already surpasses the base VLMs (LLaVA-v1.5 and LLaVA-MORE) in both hallucination reduction and content coverage. Moreover, with a lower threshold (), our model can even outperform GPT-4V in hallucination control.
| Components | CHAIR | Cover | Hall | Cog |
@@
Trade-offs Between Inference Efficiency and De-Hallucination REVERSE combines online self-verification with re-generation, introducing minimal computational overhead. To assess this, we analyze 1,004 samples from the AMBER-G test set. Verification is triggered in only 63% of cases and is highly efficient, relying on lightweight logit-based checks as opposed to a heavy pipeline in prior post-hoc verification method [51]. In the remaining 37% where verification is skipped, the CHAIR score is already low (5.1 vs. the average of 7.2), indicating that our verification can identify and skip high-quality outputs.
Among the 63% of cases where backtracking is triggered, no grounded noun phrases are mistakenly corrected to hallucinated ones. Approximately 1.5% of all cases contain hallucinations, over half of which are resolved with a single re-generation round. Furthermore, re-generation can be optimized by reusing KV-cache [43, 25], eliminating the need to recompute the full prefix. Overall, the self-correction loop in REVERSE is computationally affordable, and its effectiveness is demonstrated in Table 1, Table 2, and Table 3.
@@ Acknowledgments
Appendix C discusses the potential limitations and social impact of this work.
Appendix D describes the evaluation datasets and metrics.
Appendix E provides more qualitative and quantitative results.
Appendix F details the training and evaluation implementation for REVERSE.

### v3

@@ 1 Introduction
In this paper, we introduce REVERSE (REtrospective VERification and SElf-correction), the first framework that unifies generation, verification, and correction within a single VLM (Figure 1), enabling on-the-fly correction during decoding without waiting for the model to complete its initial output. REVERSE consists of two key novel components. First, our method fine-tunes a VLM on a specially constructed training dataset consisting of synthetic hallucination phrases tagged with a special, explicit confidence token. Unlike prior VLMs instructed with only well-grounded data, our resulting hallucination-aware model is now able to tag likely phrase-level hallucinations during the generation process.
Second, we introduce retrospective resampling, a technique that allows the hallucination-aware VLM to serve as its own verifier. During the generation process, when the hallucination-aware VLM places sufficient probability on the special hallucination token we trigger a backtracking self-correction process. Specifically, we backtrack to a previous confident section and then apply rejection sampling and query-rewriting to correct the hallucination. As illustrated in Figure 1, the introduction of explicit confidence-token training and the backtracking-based inference algorithm enable VLMs to perform interpretable and controllable self-correction, a paradigm not explored in prior work.
@@ 4 Experiments
Implementation Details We applied our method, REVERSE, on three VLM backbones: LLaVA-v1.5 (7B) [35], LLaVA-More (LLaVA with Llama-v3.1 8B) [17, 1], and Qwen2.5-VL (3B) [6]. Since LLaVA provides both its pre-trained model and instruction tuning data, we performed LoRA fine-tuning on the pre-trained model directly with our modified cross-entropy loss (see subsection 3.2) and the 1.3M-sample dataset for one epoch. In contrast, Qwen2.5-VL does not release its instruction tuning data. To enable a fair comparison, we perform full fine-tuning on the publicly available Qwen2.5-VL model using two alternatives: a 100k subset of LLaVA’s instruction data and a matched subset from our dataset. Although a more direct evaluation using Qwen2.5-VL’s original instruction tuning data and the augmentation method described in subsection 3.1 would be ideal, it is not feasible given the current release conditions. More details on training recipes are provided in Appendix E.
@@
Evaluation Protocol To evaluate our approach, we compare REVERSE with various hallucination mitigation methods, including training-free or training-based generative adjustment techniques [28, 23, 16, 3, 55, 41, 56, 58], and post-hoc verification with refinement [53]. All methods are evaluated on both VLM backbones under consistent settings, where we fix the decoding temperature at 0 and use only the base prompts provided by each dataset to ensure fair comparisons. Since REVERSE does stochastic sampling at inference time, we report the mean performance over 100 bootstrapped runs for robustness. The exact numbers with 95% confidence intervals are provided in Appendix D.
Our evaluation dataset covers several standard VQA tasks aimed at assessing visual hallucination, with a primary focus on image captioning and open-ended question answering. While discriminative tasks, binary (Yes/No) questions targeting object, attributes, and spatial understanding, are also common, backtracking provides limited benefit in such settings, which have been noted to offer less diagnostic insight into VLM hallucinations [7, 41]. We include results on these tasks in Appendix D for completeness.
For image captioning, we use CHAIR-MSCOCO [40, 55] and the generative subset of AMBER [47] (denoted as AMBER-G). CHAIR-MSCOCO evaluates object hallucination using the CHAIR score, which measures the degree of misalignment between objects mentioned in a model-generated caption and objects actually present in the image. It is defined as 1 minus the intersection over union (IoU) between the sets of mentioned and ground-truth objects. We report both CHAIRi, which aggregates the CHAIR score across all object instances, and CHAIRs, which quantifies the proportion of images where at least one hallucination occurs. For AMBER-G, we report four key metrics: CHAIR, Coverage (Cover), Hallucination (Hall), and Cognition (Cog). CHAIR is the same as CHAIRi above, and Coverage measures how well the caption mentions all objects in the image, similar to recall. The definitions of the remaining two metrics and further details on these datasets are provided in Appendix C.
@@ 4.1 Experimental Results
Open-ended Question Answering We evaluate REVERSE on MMHal-Bench and HaloQuest, two open-ended VQA benchmarks containing lots of questions with false premises or insufficient context. These examples require the model to either refuse or correct the query. For these questions, we observed that REVERSE often produces empty responses and we interpret this behavior as the model identifying the query as unanswerable. In all such cases, we apply query rewriting with the prompt: “For this question, please point out the false premises or note what information is missing, rather than answering it directly.” The complete mechanism for handling unanswerable questions is given in Appendix E.
@@ 4.2 Discussions
We conduct additional experiments on two LLaVA models to examine key aspects of our method, including ablation studies, trade-offs between expressiveness and performance, efficiency versus accuracy, and the effect of temperature. Potential limitations and broader social impacts are discussed in section 5.
Ablation Studies We conduct ablation studies to evaluate the contributions of different components of our method, as shown in subsection 4.2. Comparing the first and second rows, we observe that hallucination-aware training alone already improves performance across all metrics, outperforming existing VLMs. We hypothesize that this improvement arises from the model’s ability to contrast positive and negative phrases, effectively learning to distinguish between </CN> and </UN> during training—a mechanism that may be similar to DPO [39]. This finding suggests a potential research direction for future work. Interestingly, even a naive rejection sampling strategy reduces CHAIR hallucination scores by 1.2. When combined with query rewriting, the coverage improves by a further 1.2 points, indicating that rewriting helps the model explore alternative phrasing and correct itself more effectively.
| Components | CHAIR (↓) | Cover (↑) | Hall (↓) | Cog (↓) |
@@
| # Rounds (N) | 0 | 5 | 10 | 20 | 50 |
| CHAIR (↓) | 7.8 | 7.1 | 6.8 | 6.7 | 6.0 |
| #Tokens (%) | 1.00 | 1.75 | 2.05 | 2.63 | 3.05 |
Trade-offs Between Inference Efficiency and Hallucination To quantify the impact of REVERSE on efficiency, we analyze 1,004 samples from the AMBER-G test set. As shown in subsection 4.2, we report the number of correction rounds, hallucination scores (CHAIR), and relative runtime, measured by the total number of generated tokens of REVERSE-v1.5-7B compared to the LLaVA-v1.5-7B baseline. To show the trade-off of correction versus efficiency, these experiments are conducted under a protocol where the model continues generation after correction attempts (rather than terminating early, as our algorithm normally would). The overall overhead can be decomposed into two components: verification overhead and correction overhead. For the verification overhead, our token-level confidence estimation is performed inline during generation and introduces negligible additional cost. In contrast, prior methods like Woodpecker [53] and LURE [57] rely on auxiliary VLMs or external object detectors, incurring substantial inference overhead.
As for the correction overhead, we further evaluate performance across different maximum correction rounds. In prior work, each correction round typically entails full re-generation, leading to approximately runtime for correction rounds. In REVERSE, despite setting the maximum number of correction rounds to 50, 37% of samples require no backtracking, and among the remaining cases, over half converge within a single round. This keeps the average overhead reasonable. Additional correction rounds continue to lower CHAIR scores while runtime remains within roughly of the baseline, since corrections are localized rather than full re-generations. Moreover, the re-generation process can be further optimized by reusing KV-cache [45, 26], avoiding recomputation of the shared prefix.
Overall, the self-correction loop in REVERSE is computationally affordable and effective, as demonstrated in Table 1, Table 2, and Table 3. Yet, we believed that developing even more efficient correction strategies remains a promising avenue for future work.
Trade-offs Between Expressiveness and Hallucination As discussed in subsection 3.3, a key component of our retrospective resampling method is the predefined threshold . When the predicted probability of </UN> exceeds , the model triggers backtracking and self-correction. Figure 5 presents an analysis of the effect of on two VLMs. The 2D plot illustrates the performance trade-off between CHAIR (hallucination metric) and coverage across different threshold values. Based on our experiments, we selected as a global threshold for image captioning and other generative tasks, as it represents the peak of the performance frontier. The strong results under a single universal setting demonstrate the generalization ability of our method across diverse domains.
Moreover, the existence of threshold tuning provides an interpretable, user-adjustable control for balancing creativity and trustworthiness, which is an ability unique to our method. Compared with prior hallucination-reduction approaches, REVERSE enables dynamic adjustment of this balance between expressiveness and reliability. Notably, with a relatively high threshold (), our method already surpasses the base VLMs (LLaVA-v1.5 and LLaVA-MORE) in both hallucination reduction and content coverage. Conversely, with a lower threshold (), our model can even outperform GPT-4V in hallucination control. To our knowledge, REVERSE is the first approach to expose this trade-off transparently through a single parameter, whereas prior methods fix such thresholds implicitly without offering user control. Future work may explore how to adaptively adjust based on contextual factors or task difficulty.
@@
5 Limitations and Societal Impact
Our 1.3M Instruction Tuning Dataset
REVERSE-VLMs
REVERSE is a generalizable framework applicable to models like LLaVA-v1.5, LLaVA-MORE, and Qwen2.5-VL. It significantly reduces hallucination with minimal loss in expressiveness and efficiency (shown in Table 1, Table 2). However, it does not improve performance on discriminative VQA tasks (see Table A.1) as the backtracking methods generally didn’t help further reasoning. Future work can explore integrating REVERSE with existing hallucination reduction techniques to improve such tasks. As discussed in subsection 4.2, REVERSE reduces the hallucination with the cost of more generated tokens. An efficient self-correction methods can be a future direction.
@@ Acknowledgments
NeurIPS Paper Checklist
1.
Claims
Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
Answer: [Yes]
Justification: Claims are justified through experimental results in the paper (subsection 4.1).
Guidelines:
The answer NA means that the abstract and introduction do not include the claims made in the paper.
The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
2.
Limitations
Question: Does the paper discuss the limitations of the work performed by the authors?
Answer: [Yes]
Justification: See section 5.
Guidelines:
The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
The authors are encouraged to create a separate "Limitations" section in their paper.
The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
3.
Theory assumptions and proofs
Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
Answer: [N/A]
Justification: No theoretical results.
Guidelines:
The answer NA means that the paper does not include theoretical results.
All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
All assumptions should be clearly stated or referenced in the statement of any theorems.
The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
Theorems and Lemmas that the proof relies upon should be properly referenced.
4.
Experimental result reproducibility
Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
Answer: [Yes]
Justification: We publicly release the dataset, training code, and model checkpoints. The data curation pipeline is illustrated in Figure A.1, and training and evaluation details are provided in Appendix E.
Guidelines:
The answer NA means that the paper does not include experiments.
If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
5.
Open access to data and code
Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
Answer: [Yes]
Justification: Code, dataset, and model checkpoints will be released with the camera-ready version in Appendix A.
Guidelines:
The answer NA means that paper does not include experiments requiring code.
Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
6.
Experimental setting/details
Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
Answer: [Yes]
Justification: See section 5.
Guidelines:
The answer NA means that the paper does not include experiments.
The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
The full details can be provided either with the code, in appendix, or as supplemental material.
7.
Experiment statistical significance
Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
Answer: [Yes]
Justification: For all experiments, we compute 100 bootstrapped rounds. We only report average score in subsection 4.1 for clarity but provide detailed 95% confidence interval under all settings in Table A.5, Table A.6, and Table A.7.
Guidelines:
The answer NA means that the paper does not include experiments.
The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
The assumptions made should be given (e.g., Normally distributed errors).
It should be clear whether the error bar is the standard deviation or the standard error of the mean.
It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
8.
Experiments compute resources
Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
Answer: [Yes]
Justification: See Appendix E.
Guidelines:
The answer NA means that the paper does not include experiments.
The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
9.
Code of ethics
Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
Answer: [Yes]
Justification: We respect the code of ethics.
Guidelines:
The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
10.
Broader impacts
Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
Answer: [Yes]
Justification: See section 5.
Guidelines:
The answer NA means that there is no societal impact of the work performed.
If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
11.
Safeguards
Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
Answer: [Yes]
Justification: See section 5.
Guidelines:
The answer NA means that the paper poses no such risks.
Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
12.
Licenses for existing assets
Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
Answer: [Yes]
Justification: See Appendix A.
Guidelines:
The answer NA means that the paper does not use existing assets.
The authors should cite the original paper that produced the code package or dataset.
The authors should state which version of the asset is used and, if possible, include a URL.
The name of the license (e.g., CC-BY 4.0) should be included for each asset.
For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
13.
New assets
Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
Answer: [Yes]
Justification: Documentation is released as part of the code release.
Guidelines:
The answer NA means that the paper does not release new assets.
Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
The paper should discuss whether and how consent was obtained from people whose asset is used.
At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
14.
Crowdsourcing and research with human subjects
Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
Answer: [N/A]
Justification: N/A
Guidelines:
The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
15.
Institutional review board (IRB) approvals or equivalent for research with human subjects
Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
Answer: [N/A]
Justification: N/A
Guidelines:
The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
16.
Declaration of LLM usage
Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
Answer: [Yes]
Justification: GitHub Copilot was used during coding. In paper writing, LLM was used solely for grammar checking.
Guidelines:
The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.
@@
Appendix C describes the evaluation datasets and metrics.
Appendix D provides more qualitative and quantitative results.
Appendix E details the training and evaluation implementation for REVERSE.